hhuuc:

(3.14 x 10^23 ly) - (7.18 x 10^20 ly) Need Help Solving This..

jhonyy9:

- so for you can making this substraction you need rewriting the first expression in a form with 10^20 ,like in case of the second expression

hhuuc:

ok

jhonyy9:

hope helped

hhuuc:

yeah kinda, thanks for responding

jhonyy9:

can you solve it now sure ?

hhuuc:

ill try im not that good in physics tho

jhonyy9:

can you writing here now for i can check it - ok. ?

jhonyy9:

3,14×10^23 = ? ×10^20 ?

hhuuc:

3.14e + 43 Thas what i got

jhonyy9:

i dont know how you ve got this continuing my above wrote idea 3,14×10^23 = 3140×10^20 because take the decimal 3 units on the right - yes ?

hhuuc:

ohh

jhonyy9:

and now there are 3140,00×10^20 - 7,18×10^20

jhonyy9:

can you making this substraction now ?

hhuuc:

yeah thanks

jhonyy9:

i wait it

jhonyy9:

please write here

hhuuc:

3.13282+23

jhonyy9:

now bc. the ten power parts are the same so you need using the left part of 3140 - 7,18 = 3132,82 but there is again the ten part with exponent 20 so will be 3132,82×10^20 ly

hhuuc:

Damn i have a test about this tomorrow, guess i am going to get a bad grade

hhuuc:

i gotta do 7 more exercises about this..

jhonyy9:

why ? your solution is right - what you ve wrote above looke please than we need using power of 10 of 23 so from 3132,82×10^20 we need making expression with 10 on power of 23 yes ? so just take the decimal now on the left part 3 units and will get the result what you ve wrote above 3132,82×10^20 = 3,13282×10^23 hope you understand it now sure

jhonyy9:

this result is exactly what you ve wrote above

hhuuc:

kk

jhonyy9:

@SmokeyBrown

hhuuc:

well thanks for the help bro, Have a good one..

jhonyy9:

np. was my pleasure and any time bye bye but do you understand it now ?

hhuuc:

yeah i just need more practice thats all

jhonyy9:

exactly - you see it right - good work and sucesses